Write the complex number in the form of a+bi? 5/2(cos150degrees+i sin150degrees)
1
Expert's answer
2012-03-28T16:16:12-0400
cos(150)=-cos(30)=-sqrt(3)/2
sin(150)=sin(30)=1/2
So we get
5/2(cos150degrees+i sin150degrees)=5/2*(-sqrt(3)/2+i/2)=-5sqrt(3)/4+5i/4